OpenStudy (anonymous):

will give medal Which statements are true according to the order of operations? Choose exactly two answers that are correct. A. 32 + 8 ÷ 4 = 40 ÷ 4 = 10 B. 15 • (6 ÷ 3) = 90 ÷ 3 = 30 C. 4 + 12 ÷ 2 = 4 + 6 = 10 D. (20 – 5) • 3 = 15 • 3 = 45

OpenStudy (haichi):

C and D

OpenStudy (haichi):

12/2=6 and 6 +4=10

OpenStudy (haichi):

20-5=15 15*3=45